Evaluate 12!/5!7! WITH OUT A CALCULATOR PLEASE and show complete answer thanks!
Found 3 solutions by Alan3354, MathLover1, MathTherapy:
Answer by Alan3354(69443)

Evaluate 12!/5!7!
--------
= 12*11*10*9*8/5*4*3*2*1
= 11*10*9*8/5*2*1
= 11*9*8/1
= 792

Answer by MathLover1(20850)
Answer by MathTherapy(10552)

Evaluate 12!/5!7! WITH OUT A CALCULATOR PLEASE and show complete answer thanks!